# Renormalisation for Reaction-Diffusion Systems with Non-Local Interactions

### v1

@@ 1 Introduction
Reaction-diffusion processes are an enveloping term describing systems with discrete, mobile interacting entities [Liggett1985]. These can be used to describe systems on many scales, ranging from molecular reactions [Doi1976, Doi1976b] to spatial predator-prey systems [Dobramysl2018], for example. Although the objects in such populations differ substantially in nature, they shall be referred to as particles from this point forward. Techniques to investigate such models can be quite varied, including semigroup approaches [Banasiak2006, Goldstein2017], stochastic approaches [Gardiner1985, Liggett1985], sample size methods [Van1992], and field theory techniques [Tauber2014] to name a few.
@@
It proves useful to describe the interactions algebraically. A reaction written as $A+A\rightarrow\phi$, for example, would indicate either pairs of colliding particles that annihilate, or pairs that combine into a different class, say $B$, which are not of interest and subsequently ignored. This is a paradigm model for such systems [Doi1976, Doi1976b, Lee1994]. This also has a natural extension where branching is also included $A\rightarrow A+A$ [Cardy1998]. Spatial effects with some form of diffusion is usually implicit with these models. Such processes can contain critical points of phase transition. In particular, for pure annihilation, the asymptotic large time behaviour exhibits a qualitative change at critical spatial dimension $d_{c}=2$. If branching is added to the process, a different critical behaviour emerges; the point at which branching and annihilation rates conspire to produce a non-zero steady state has critical dimension $d_{c}=4$. Renormalisation methods of field theory have proven useful for examining criticality of such systems [Lee1994, Peliti1986, Tauber2014] and is the subject of this study.
@@
Usually, these methods are developed with the aid of path integral techniques. These often start with an underlying lattice model [Lee1994, Peliti1985] such that pairs of particles can only annihilate or spawn on single lattice sites. Such systems can then limit to a continuum process, with a ‘lattice constant’ $a$ (indicating separation between proximal sites) being reduced to zero. However, for many systems of interest, either because of particle size, or interactions acting over a distance, such local models are insufficient to capture this behaviour. For example, the original work of Doi [Doi1976, Doi1976b] considered non-local interactions (albeit without path integrals) for the annihilation model $A_{p}+A_{q}\rightarrow\phi$, where particles at positions $p$ and $q$ annihilate when sufficiently close to each other. If $p$ instead represents the age of a particle (rather than position), a birth process can be written as $A_{p}\rightarrow A_{p}+A_{0}$, where a new offspring with age zero appears. This is inherently non-local [Greenman2017], and is also an example of stochastic resetting which has recently generated significant interest [Evans2020]. Dealing with non-local interactions has also been considered in the guise of polymer reactions [DeGennes1982, Kleinert2006, Kleinert2006b]. In some cases the polymer is a string and interactions occur during contact (so in some sense is still local), but in some cases, the polymers have width and the interactions are non-local. Adapting renormalisation methods to deal with such non-local behaviour is thus of interest and the primary aim of this work.
@@
Utilising renormalisation techniques to capture the asymptotic behaviour of the systems generally encompasses two features [Lee1994, Tauber2014]. Firstly, perturbation expansions in field theory often involve ‘ultra violet’ (UV) divergent terms, which renormalisation methods can systematically remove (albeit just for renormalisable theories). Secondly, the asymptotic behaviour of interest often involve ‘infra red’ (IR) divergences in critical regions of interest, and renormalisation group techniques are used to connect features in these divergent domains near critical points to perturbative terms arising from convergent regions.
@@
In this work, path integral renormalisation methods for non-local interactions are considered in detail. Specifically, we show that non-local interactions have the natural effect of regulating UV divergences, which either have reduced divergence, or no longer diverge below the critical dimension of interest. However, the IR divergences still exist and need to be treated to examine asymptotics. The non-local nature of interactions can make the perturbative re-summations needed for renormalisation somewhat intractable. However, the rescaling that underlies the renormalisation group will be seen to have the convenient effect of converting non-local into local interactions. Subsequently, the critical dimensions and associated universal properties observed for locally interacting systems, that demarcate a dimensional transition in asymptotic decay exponents, are preserved. We show this for both the annihilation process [Lee1994] and when branching, spontaneous birth and death is included [Cardy1998].
@@ 2 Field Theory
This section contains the following. Firstly, some conventions are specified. Then, to fully define the models of interest, the non-local nature of some interactions are specified. Doi-Peliti methods for constructing path integrals are then introduced. This is done without recourse to lattice methods and is more akin to Doi’s original continuous approach [Doi1976, Doi1976b]. The non-local nature of interaction leads to path integral actions that have double integrals over space, which are awkward to deal with. A Hubbard-Stratonovich transform is then shown to disentangle this, at the price of two extra fields. Investigation of mean field behaviour completes this section, providing the necessary background for subsequent renormalisation methods.
@@ 2.2 Interaction Profiles
Before describing some possible interaction profiles, we briefly summarise the different effects that local and non-local models of interaction have on Feynman diagrams. In Fig. 1A we see a divergent diagram arising in the pure annihilation model of [Lee1994] with local interaction. This diagram has two constant interaction terms $R$, and two propagators, each with a quadratic denominator in momentum space (see [Lee1994] or Eq. (6) for details). Integration over momentum and frequency (dimension $d+2$) thus gives a superficial degree of divergence of d2d-2. Once non-local effects are enabled, the interaction constants $R$ becomes functions $R(k)$ in momentum space (Fig. 1B). Depending on the form of interaction (see below), these conspire to further reduce the degree of divergence and diagrams are less divergent for non-local interactions. For example, if $R(k)$ has a quadratic denominator, the degree of divergence becomes d6d-6, whereas if $R(k)$ decays exponentially, the diagram is convergent in any dimension. Both cases arise for specific examples of interest described below.
@@
Figure 1: A) Single loop Feynman diagram for the annihilation model in [Lee1994] with local interaction rate $R$. B) Corresponding diagram for non-local interaction function $R(k)$ (in momentum representation). Dotted lines indicate truncated propagators. In both cases external momenta are zero and the loop momentum is $k$.
@@
Normal $R\left(\frac{\lambda}{\pi}\right)^{\frac{d}{2}}e^{-\lambda r^{2}}$ $Re^{-\frac{k^{2}}{4\lambda^{2}}}$ $\frac{1}{2\lambda}$
Screened Poisson $R(2\pi)^{-\frac{d}{2}}\left(\frac{\lambda}{r}\right)^{\frac{d-2}{2}}K_{\frac{d-2}{2}}\left(\lambda r\right)$ $R\frac{\lambda^{2}}{\lambda^{2}+k^{2}}$ $\frac{2}{\lambda^{2}}$
Spherical $\frac{R}{V_{d}}\theta(|r|<\frac{1}{\lambda})$ $R(\frac{2\lambda}{k})^{\frac{d}{2}}\Gamma(\frac{d}{2}+1)J_{\frac{d}{2}}(\frac{k}{\lambda})$ $\frac{1}{2\lambda^{2}}\frac{\Gamma(\frac{d}{2}+\frac{1}{2})}{\Gamma(\frac{d}{2}+2)}$
@@
Table 1: A range of interaction profiles $R(r)=R\hat{R}(r)$ that are a function of separation $r$. The parameter $\lambda$ in all cases represents a form of precision. The function $J_{\frac{d}{2}}$ is a Bessel function of the first kind. The function $K_{\frac{d-2}{2}}$ is a modified Bessel function of the second kind. The term $V_{d}=\frac{\pi^{\frac{d}{2}}}{\Gamma(\frac{d}{2}+1)}$ is the volume of a $d$ dimensional unit sphere. The Reisz potential is defined provided the test function $f$ decays sufficiently quickly. The value $c_{\nu}=\pi^{\frac{d}{2}}2^{\nu}\frac{\Gamma(\frac{\nu}{2})}{\Gamma(\frac{d-\nu}{2})}$ is a Fourier transform normalisation term, and $N_{f,\nu}$ is a normalisation constant to ensure a total integral of $R$.
@@
The Local form $R(r)=R\delta(r)$ is the profile used for most lattice models [Lee1994, Peliti1985], with interactions occurring on individual lattice sites. Note that this takes the constant form $R(k)=R$ in momentum space, contributing no terms of consequence when ultra-violet divergence is considered. All the remaining cases considered are non-local, with a form $R(r)=R\hat{R}(r)$ involving an overall rate $R$ multiplied by a normalised probability density $\hat{R}(r)$ that is a function of separation, $r$ (apart from the Reisz potential described below). The Normal case gives a natural form of interaction, with a strongly decaying function in momentum space, which will be seen to provide a strong form of regulation for UV divergence. The Screened Poisson distribution generalizes the Coulomb potential into a normalizable function, satisfying an equation of the form $(\nabla^{2}-\lambda^{2})\hat{R}=-\delta(r)$. This structure also has a convenient Fourier transform $\frac{\lambda^{2}}{\lambda^{2}+k^{2}}$ which will prove pliable when Feynman integrals in momentum space are considered. Note the quadratic nature of the denominator, which will be seen to provide some regulation when ultra-violet divergences are considered. The Spherical profile is perhaps the most natural. This can be interpreted as giving particles a finite, specific width, with interactions occurring at fixed rates when particles touch. It can also be interpreted as interactions between point particles that require a proximity threshold. This is the form considered with field theory by Doi [Doi1976, Doi1976b], originally considered as interaction on contact by Smoluchowski [Von1917] and later as an interaction rate by Teramoto and Shigesada [Teramoto1967]. Although the spherical model is a natural choice, the hard edges of the interaction profile do not produce a very forgiving function in momentum space. However, asymptotically, this takes the form $\mathcal{O}\left({k^{-\frac{d+1}{2}}}\right)$ and so provides weaker regulation than the screened Poisson for $d\simeq 2$. Finally, the Reisz potential can provide varied regulation, in particular, it can provide regulation between the delta function (none) and the screened Poisson (quadratic), although analytic results will only be possible for certain choices of test function $f$.
@@ 2.3 Doi-Peliti Methods
Here the requisite Doi-Peliti field theory is developed. This material can be found elsewhere and is just an overview. Note that the original development by Doi [Doi1976, Doi1976b] does not use a discrete lattice, but is entirely continuous in nature. It does not treat path integrals, however, which are needed. Many of the path integral formulations seen today [Peliti1985, Tauber2014] utilise lattice formulations. The description below does not, and directly implements a continuous path integral framework [Greenman2017]. Although this offers an alternative derivation and a point of interest, being more closely aligned with Doi’s original approach, the end results are the same as those of lattic methods.
@@
A slightly different formulation is used in [Greenman2017], which results in path integrals over spaces of real functions, rather than two conjugate complex fields.
To build a stochastic model, we start with a suitable Liouvillian operator to capture the dynamics of interest. To do this the associated models need to be specified. The first, Model I, is a pure annihilation model, and is a non-local version of the model considered by Lee [Lee1994]. The second model generalizes this, and includes processes of branching, spontaneous birth and death, giving a non-local version of the model considered by Janssen [Janssen1999]. Both models assume an underlying diffusion at rate $D$, and are summarized below:
@@
where the last term utilises the resolution of identity to form a path integral via the usual methods of time slicing [Greenman2017, Peskin1995, Sredniki2007, Tauber2014], which converts the Liouvillian operator into an action $S$. The normalisation constant $\mathcal{N}=\int\mathcal{D}[\psi_{p,t}]\mathcal{D}[\bar{\psi}_{p,t}]e^{S}=1$ is just the total probability density and could be ignored at this point. However, Jacobians in corresponding path integral measures will appear during scaling transformations, which will subsequently cancel if the term $\mathcal{N}$ is included. Implementing a standard Doi field shift $\bar{\psi}_{p}\rightarrow\bar{\psi}_{p}+1$ finally results in the following action (for the full Model II),
@@
The pure annihilation Model I just retains the first line in this equation. Note that the time dependences of the fields are now suppressed for simplicity of notation (the parameters are taken to be time independent). There is also an important simplification here. Specifically, we assume that the initial spread of particles is uniform. That is, we take initial density $\omega(p)=\frac{1}{V}$ per particle over some large volume $V$. As the mean initial number of particles is Poisson with mean value $m$, the corresponding term in the action then becomes $m\int_{p}\omega_{p}\bar{\psi}_{p}(t=0)=\frac{m}{V}\int_{p}\bar{\psi}_{p}(t=0)$. Defining the value $n_{0}=\frac{m}{V}$ as the particle density then reproduces the corresponding term in the action above. Note that if we implement the local form of interaction $R_{pq}=R\delta(p-q)$, remove branching ($Q_{pq}=0$), spontaneous birth ($B_{p}=0$) and death ($M_{p}=0$), then the site specific local annihilation action seen in [Lee1994] is recovered.
@@ 2.4 Hubbard-Stratonovich Transform
Expansions of the action can be simplified with the aid of a Hubbard-Stratonovich transform. These transforms basically introduce two auxiliary fields into the action, which allows the separation of complicated terms into simpler components [Coleman1988, Hofmann2022, Kopietz1997]. For the situation in hand, specifically, the following identity is required,
@@
For the case ($M=0$), this reduces to the standard propagator found in annihilation models [Lee1994]. The other two quadratic terms result in propagators (in momentum space) $R(k)$ and $Q(k)$. The lack of time dependence (for interaction rates) means that the start and end of propagators have the same time point (and so are drawn vertically in Fig. 2). Note that all interactions we consider are isotropic, meaning in particular that R(x)=R(x)R(x)=R(-x), and so in Fourier space, R(k)=R(k)R(k)=R(-k), meaning the coefficient of the undirected edge in Fig. 2C can be taken as simply R(k)-R(k). The resulting Feynman graph components are summarised in Fig. 3B.
Note, firstly, that directions on edges are clear from context and can be dropped. Secondly, the remaining terms, dealt with by perturbative expansion, now only implicate a single position variable. Thirdly, interaction functions and parameters are now only present as propagators, meaning the nodes only have numerical coefficients, except for terminal nodes containing a $B(k)$ term (see Fig. 2F). Finally, the rules of expansion restricted to annihilation match those found in [Doi1976, Doi1976b], where time ordering (that is, not path integral) formulation of field theory is used. They can also be obtained utilising quantum stochastic process machinery [Greenman2023].
@@
Note that the local case $Q(k)=DQ$ recovers the annihilation-branching propagator found in [Cardy1998].

### v2

@@ 1 Introduction
Reaction-diffusion processes are an enveloping term describing systems with discrete, mobile interacting entities [24]. These can be used to describe systems on many scales, ranging from molecular reactions [10, 11] to spatial predator-prey systems [9], for example. Although the objects in such populations differ substantially in nature, they shall be referred to as particles from this point forward. Techniques to investigate such models can be quite varied, including semigroup approaches [1, 14], stochastic approaches [13, 24], sample size methods [34], and field theory techniques [32] to name a few.
@@
It proves useful to describe the interactions algebraically. A reaction written as $A+A\rightarrow\phi$, for example, would indicate either pairs of colliding particles that annihilate, or pairs that combine into a different class, say $B$, which are not of interest and subsequently ignored. This is a paradigm model for such systems [10, 11, 23]. This also has a natural extension where branching is also included $A\rightarrow A+A$ [5]. Spatial effects with some form of diffusion is usually implicit with these models. Such processes can contain critical points of phase transition. In particular, for pure annihilation, the asymptotic large time behaviour exhibits a qualitative change at critical spatial dimension $d_{c}=2$. If branching is added to the process, a different critical behaviour emerges; the point at which branching and annihilation rates conspire to produce a non-zero steady state has critical dimension $d_{c}=4$. Renormalisation methods of field theory have proven useful for examining criticality of such systems [23, 26, 32] and is the subject of this study.
@@
Usually, these methods are developed with the aid of path integral techniques. These often start with an underlying lattice model [23, 25] such that pairs of particles can only annihilate or spawn on single lattice sites. Such systems can then limit to a continuum process, with a ‘lattice constant’ $a$ (indicating separation between proximal sites) being reduced to zero. However, for many systems of interest, either because of particle size, or interactions acting over a distance, such local models are insufficient to capture this behaviour. For example, the original work of Doi [10, 11] considered non-local interactions (albeit without path integrals) for the annihilation model $A_{p}+A_{q}\rightarrow\phi$, where particles at positions $p$ and $q$ annihilate when sufficiently close to each other. If $p$ instead represents the age of a particle (rather than position), a birth process can be written as $A_{p}\rightarrow A_{p}+A_{0}$, where a new offspring with age zero appears. This is inherently non-local [15], and is also an example of stochastic resetting which has recently generated significant interest [12]. Dealing with non-local interactions has also been considered in the guise of polymer reactions [8, 21, 20]. In some cases the polymer is a string and interactions occur during contact (so in some sense is still local), but in some cases, the polymers have width and the interactions are non-local. Adapting renormalisation methods to deal with such non-local behaviour is thus of interest and the primary aim of this work.
@@
Utilising renormalisation techniques to capture the asymptotic behaviour of the systems generally encompasses two features [23, 32]. Firstly, perturbation expansions in field theory often involve ‘ultra violet’ (UV) divergent terms, which renormalisation methods can systematically remove (albeit just for renormalisable theories). Secondly, the asymptotic behaviour of interest often involve ‘infra red’ (IR) divergences in critical regions of interest, and renormalisation group techniques are used to connect features in these divergent domains near critical points to perturbative terms arising from convergent regions.
@@
In this work, path integral renormalisation methods for non-local interactions are considered in detail. Specifically, we show that non-local interactions have the natural effect of regulating UV divergences, which either have reduced divergence, or no longer diverge below the critical dimension of interest. However, the IR divergences still exist and need to be treated to examine asymptotics. The non-local nature of interactions can make the perturbative re-summations needed for renormalisation somewhat intractable. However, the rescaling that underlies the renormalisation group will be seen to have the convenient effect of converting non-local into local interactions. Subsequently, the critical dimensions and associated universal properties observed for locally interacting systems, that demarcate a dimensional transition in asymptotic decay exponents, are preserved. We show this for both the annihilation process [23] and when branching, spontaneous birth and death is included [5].
@@ 2 Field Theory
This section contains the following. Firstly, some conventions are specified. Then, to fully define the models of interest, the non-local nature of some interactions are specified. Doi-Peliti methods for constructing path integrals are then introduced. This is done without recourse to lattice methods and is more akin to Doi’s original continuous approach [10, 11]. The non-local nature of interaction leads to path integral actions that have double integrals over space, which are awkward to deal with. A Hubbard-Stratonovich transform is then shown to disentangle this, at the price of two extra fields. Investigation of mean field behaviour completes this section, providing the necessary background for subsequent renormalisation methods.
@@ 2.2 Interaction Profiles
Before describing some possible interaction profiles, we briefly summarise the different effects that local and non-local models of interaction have on Feynman diagrams. In Fig. 1A we see a divergent diagram arising in the pure annihilation model of [23] with local interaction. This diagram has two constant interaction terms $R$, and two propagators, each with a quadratic denominator in momentum space (see [23] or Eq. (6) for details). Integration over momentum and frequency (dimension $d+2$) thus gives a superficial degree of divergence of d2d-2. Once non-local effects are enabled, the interaction constants $R$ becomes functions $R(k)$ in momentum space (Fig. 1B). Depending on the form of interaction (see below), these conspire to further reduce the degree of divergence and diagrams are less divergent for non-local interactions. For example, if $R(k)$ has a quadratic denominator, the degree of divergence becomes d6d-6, whereas if $R(k)$ decays exponentially, the diagram is convergent in any dimension. Both cases arise for specific examples of interest described below.
@@
Figure 1: A) Single loop Feynman diagram for the annihilation model in [23] with local interaction rate $R$. B) Corresponding diagram for non-local interaction function $R(k)$ (in momentum representation). Dotted lines indicate truncated propagators. In both cases external momenta are zero and the loop momentum is $k$.
@@
Normal $R\left(\frac{\lambda^{2}}{\pi}\right)^{\frac{d}{2}}e^{-\lambda^{2}r^{2}}$ $Re^{-\frac{k^{2}}{4\lambda^{2}}}$ $\frac{1}{2\lambda^{2}}$
Screened Poisson $R(2\pi)^{-\frac{d}{2}}\lambda^{2}\left(\frac{\lambda}{r}\right)^{\frac{d}{2}-1}K_{\frac{d}{2}-1}\left(\lambda r\right)$ $R\frac{\lambda^{2}}{\lambda^{2}+k^{2}}$ $\frac{2}{\lambda^{2}}$
Spherical $\frac{R\lambda^{d}}{V_{d}}\theta(|r|<\frac{1}{\lambda})$ $R(\frac{2\lambda}{k})^{\frac{d}{2}}\Gamma(\frac{d}{2}+1)J_{\frac{d}{2}}(\frac{k}{\lambda})$ $\frac{1}{\lambda^{2}(d+2)}$
@@
Table 1: A range of interaction profiles $R(r)=R\hat{R}(r)$ that are a function of separation $r$. The parameter $\lambda$ in all cases represents a form of precision. The function $J_{\frac{d}{2}}$ is a Bessel function of the first kind. The function $K_{\frac{d}{2}-1}$ is a modified Bessel function of the second kind. The term $V_{d}=\frac{\pi^{\frac{d}{2}}}{\Gamma(\frac{d}{2}+1)}$ is the volume of a $d$ dimensional unit sphere. The Reisz potential is defined provided the test function $f$ decays sufficiently quickly. The value $c_{\nu}=\pi^{\frac{d}{2}}2^{\nu}\frac{\Gamma(\frac{\nu}{2})}{\Gamma(\frac{d-\nu}{2})}$ is a Fourier transform normalisation term, and $N_{f,\nu}$ is a normalisation constant to ensure a total integral of $R$.
@@
The Local form $R(r)=R\delta(r)$ is the profile used for most lattice models [23, 25], with interactions occurring on individual lattice sites. Note that this takes the constant form $R(k)=R$ in momentum space, contributing no terms of consequence when ultra-violet divergence is considered. All the remaining cases considered are non-local, with a form $R(r)=R\hat{R}(r)$ involving an overall rate $R$ multiplied by a normalised probability density $\hat{R}(r)$ that is a function of separation, $r$ (apart from the Reisz potential described below). The Normal case gives a natural form of interaction, with a strongly decaying function in momentum space, which will be seen to provide a strong form of regulation for UV divergence. The Screened Poisson distribution generalizes the Coulomb potential into a normalizable function, satisfying an equation of the form $(\nabla^{2}-\lambda^{2})\hat{R}=-\delta(r)$. This structure also has a convenient Fourier transform $\frac{\lambda^{2}}{\lambda^{2}+k^{2}}$ which will prove pliable when Feynman integrals in momentum space are considered. Note the quadratic nature of the denominator, which will be seen to provide some regulation when ultra-violet divergences are considered. The Spherical profile is perhaps the most natural. This can be interpreted as giving particles a finite, specific width, with interactions occurring at fixed rates when particles touch. It can also be interpreted as interactions between point particles that require a proximity threshold. This is the form considered with field theory by Doi [10, 11], originally considered as interaction on contact by Smoluchowski [36] and later as an interaction rate by Teramoto and Shigesada [33]. Although the spherical model is a natural choice, the hard edges of the interaction profile do not produce a very forgiving function in momentum space. However, asymptotically, this takes the form $\mathcal{O}\left({k^{-\frac{d+1}{2}}}\right)$ and so provides weaker regulation than the screened Poisson for $d\simeq 2$. Finally, the Reisz potential can provide varied regulation, in particular, it can provide regulation between the delta function (none) and the screened Poisson (quadratic), although analytic results will only be possible for certain choices of test function $f$.
@@ 2.3 Doi-Peliti Methods
Here the requisite Doi-Peliti field theory is developed. This material can be found elsewhere and is just an overview. Note that the original development by Doi [10, 11] does not use a discrete lattice, but is entirely continuous in nature. It does not treat path integrals, however, which are needed. Many of the path integral formulations seen today [25, 32] utilise lattice formulations. The description below does not, and directly implements a continuous path integral framework [15]. Although this offers an alternative derivation and a point of interest, being more closely aligned with Doi’s original approach, the end results are the same as those of lattic methods.
@@
A slightly different formulation is used in [15], which results in path integrals over spaces of real functions, rather than two conjugate complex fields.
To build a stochastic model, we start with a suitable Liouvillian operator to capture the dynamics of interest. To do this the associated models need to be specified. The first, Model I, is a pure annihilation model, and is a non-local version of the model considered by Lee [23]. The second model generalizes this, and includes processes of branching, spontaneous birth and death, giving a non-local version of the model considered by Janssen [19]. Both models assume an underlying diffusion at rate $D$, and are summarized below:
@@
where the last term utilises the resolution of identity to form a path integral via the usual methods of time slicing [15, 27, 29, 32], which converts the Liouvillian operator into an action $S$. The normalisation constant $\mathcal{N}=\int\mathcal{D}[\psi_{p,t}]\mathcal{D}[\bar{\psi}_{p,t}]e^{S}=1$ is just the total probability density and could be ignored at this point. However, Jacobians in corresponding path integral measures will appear during scaling transformations, which will subsequently cancel if the term $\mathcal{N}$ is included. Implementing a standard Doi field shift $\bar{\psi}_{p}\rightarrow\bar{\psi}_{p}+1$ finally results in the following action (for the full Model II),
@@
The pure annihilation Model I just retains the first line in this equation. Note that the time dependences of the fields are now suppressed for simplicity of notation (the parameters are taken to be time independent). There is also an important simplification here. Specifically, we assume that the initial spread of particles is uniform. That is, we take initial density $\omega(p)=\frac{1}{V}$ per particle over some large volume $V$. As the mean initial number of particles is Poisson with mean value $m$, the corresponding term in the action then becomes $m\int_{p}\omega_{p}\bar{\psi}_{p}(t=0)=\frac{m}{V}\int_{p}\bar{\psi}_{p}(t=0)$. Defining the value $n_{0}=\frac{m}{V}$ as the particle density then reproduces the corresponding term in the action above. Note that if we implement the local form of interaction $R_{pq}=R\delta(p-q)$, remove branching ($Q_{pq}=0$), spontaneous birth ($B_{p}=0$) and death ($M_{p}=0$), then the site specific local annihilation action seen in [23] is recovered.
@@ 2.4 Hubbard-Stratonovich Transform
Expansions of the action can be simplified with the aid of a Hubbard-Stratonovich transform. These transforms basically introduce two auxiliary fields into the action, which allows the separation of complicated terms into simpler components [6, 17, 22]. For the situation in hand, specifically, the following identity is required,
@@
For the case ($M=0$), this reduces to the standard propagator found in annihilation models [23]. The other two quadratic terms result in propagators (in momentum space) $R(k)$ and $Q(k)$. The lack of time dependence (for interaction rates) means that the start and end of propagators have the same time point (and so are drawn vertically in Fig. 2). Note that all interactions we consider are isotropic, meaning in particular that R(x)=R(x)R(x)=R(-x), and so in Fourier space, R(k)=R(k)R(k)=R(-k), meaning the coefficient of the undirected edge in Fig. 2C can be taken as simply R(k)-R(k). The resulting Feynman graph components are summarised in Fig. 3B.
Note, firstly, that directions on edges are clear from context and can be dropped. Secondly, the remaining terms, dealt with by perturbative expansion, now only implicate a single position variable. Thirdly, interaction functions and parameters are now only present as propagators, meaning the nodes only have numerical coefficients, except for terminal nodes containing a $B(k)$ term (see Fig. 2F). Finally, the rules of expansion restricted to annihilation match those found in [10, 11], where time ordering (that is, not path integral) formulation of field theory is used. They can also be obtained utilising quantum stochastic process machinery [16].
@@
Note that the local case $Q(k)=DQ$ recovers the annihilation-branching propagator found in [5].